OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 6 kilometers, what is the height of the cone? Use 3.14 for pi . Show your work and round to the nearest tenth.

razor99 (razor99):

hehehe

razor99 (razor99):

out of reef and out of luck buddy

OpenStudy (anonymous):

sigh* i guess ure rite. u dont get a medal though

razor99 (razor99):

i can help ya buddy i was just joking

OpenStudy (anonymous):

oh ok

razor99 (razor99):

do u know the formulae

OpenStudy (anonymous):

i know the formula for the volume of a cone

OpenStudy (anonymous):

\[r^{2}hpi/3 \]

OpenStudy (anonymous):

(1/3) pi r^2 h = 300 => h = 900 / (3.14 * 36) = 8.0

razor99 (razor99):

spinner

OpenStudy (anonymous):

wow that was unexpected

razor99 (razor99):

actually its correct bro

OpenStudy (anonymous):

@JackJones don't look at my answer, before you follow razor99's instruction!

OpenStudy (anonymous):

to late

razor99 (razor99):

brb have to help a girl bye

OpenStudy (anonymous):

bye